# GlycanML: A Multi-Task and Multi-Structure Benchmark for Glycan Machine Learning

### v2

@@ Abstract
Glycans are basic biomolecules and perform essential functions within living organisms. The rapid increase of functional glycan data provides a good opportunity for machine learning solutions to glycan understanding. However, there still lacks a standard machine learning benchmark for glycan function prediction. In this work, we fill this blank by building a comprehensive benchmark for Glycan Machine Learning (GlycanML). The GlycanML benchmark consists of diverse types of tasks including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. Glycans can be represented by both sequences and graphs in GlycanML, which enables us to extensively evaluate sequence-based models and graph neural networks (GNNs) on benchmark tasks. Furthermore, by concurrently performing eight glycan taxonomy prediction tasks, we introduce the GlycanML-MTL testbed for multi-task learning (MTL) algorithms. Also, we evaluate how taxonomy prediction can boost other three function prediction tasks by MTL. Experimental results show the superiority of modeling glycans with multi-relational GNNs, and suitable MTL methods can further boost model performance. We provide all datasets and source codes at https://github.com/GlycanML/GlycanML and maintain a leaderboard at https://GlycanML.github.io/project.
@@ 1 Introduction
There are some existing works that employ machine learning models to predict glycan taxonomy [11, 12], glycan immunogenicity [11, 12], glycosylation [42, 36] and protein-glycan interaction [40]. These works aim to solve one or several related glycan function prediction problems with either sequence models adapted from natural language processing (NLP) [11, 42] or graph neural networks (GNNs) [12, 40]. However, there still lacks a comprehensive benchmark studying the general effectiveness of various machine learning models on understanding diverse glycan functions, which hinders the progress of machine learning for glycan understanding. As a matter of fact, benchmark datasets greatly facilitate the machine learning research of other biomolecules like small molecules [27, 21] and proteins [46, 63].
@@
Therefore, in this work, we take the initiative of building a Glycan Machine Learning (GlycanML) benchmark featured with diverse types of tasks and multiple glycan representation structures. The GlycanML benchmark consists of 11 benchmark tasks for understanding important glycan properties, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. For each task, we carefully split the benchmark dataset to evaluate the generalization ability of machine learning models in real-world scenarios. For example, in glycan taxonomy prediction, we leave out the glycans with unseen structural motifs during training for validation and test, which simulates the classification of newly discovered glycans in nature with novel molecular structures.
@@
The GlycanML benchmark accommodates two glycan representation structures, i.e., glycan tokenized sequences and glycan planar graphs. For each structure, we adopt suitable machine learning models for representation learning, where sequence encoders such as CNN [24], LSTM [26] and Transformer [56] are employed for glycan sequence encoding, and both homogeneous GNNs [31, 57, 62] and heterogeneous GNNs [22, 48, 55] are used to encode glycan graphs. In addition, two state-of-the-art small molecule encoders [66, 45] are also studied for all-atom glycan modeling. We evaluate each model on all benchmark tasks to study its general effectiveness. The GlycanML benchmark also provides a testbed, namely GlycanML-MTL, for multi-task learning (MTL) algorithms, where an MTL method is asked to simultaneously solve eight glycan taxonomy prediction problems which are highly correlated. The performance on this testbed measures how well an MTL method can transfer the knowledge learned from different glycan taxonomies, e.g., transferring between species-level classification and genus-level classification. Taking a step further, we study the effect of MTL on boosting three function prediction tasks (i.e., immunogenicity, glycosylation and interaction prediction) by concurrently performing taxonomy prediction.
@@
Benchmark results show that the RGCN model [48], a typical heterogeneous GNN, performs best on most benchmark tasks, and a simple two-layer CNN can surprisingly achieve competitive performance by using only condensed sequential information of glycan structures. The MTL methods with elaborate task-reweighting strategies can further enhance the performance of glycan taxonomy prediction, and learning the immunogenicity and glycosylation type of glycans jointly with their taxonomies also achieves benefits, showing the potential of MTL for glycan understanding. We hope the GlycanML benchmark will spark the interest of studying glycoscience with machine learning.
@@ 2 Related Work
Glycan machine learning. With the expanding size of experimental glycomics datasets, the integration of machine learning techniques into glycoinformatics shows considerable promise [8, 36]. Early approaches that adapt machine learning into glycomics research include the use of traditional ML algorithms (e.g., SVMs) to learn patterns from mass spectrometry data [33, 37], predict glycosylation sites [13, 35, 51, 44], and classify glycans [64]. Recently, thanks to the advancements in deep learning and new glycomics datasets, there has been a rise in studies applying deep learning to glycan and glycosylation modeling. DeepNGlyPred [42] seeks to identify N-glycosylated sequon from the N-GlyDE dataset [44]. SugarBase [10] is a comprehensive glycan database with metadata and analytical tools, and it supports the development of many glycan representation learning models, such as SweetOrigins [10], SweetTalk [9], SweetNet [12], glyBERT [18] and GNNGLY [1]. Deep learning-based methods like LectinOracle [40] and GlyNet [14] also made notable progress in predicting binding strengths between proteins and glycans.
@@
Biological machine learning benchmarks. To evaluate the performance of different machine learning methods in modeling biomolecules, it is necessary to establish large-scale standardized benchmarks. MoleculeNet [61] is a widely-used benchmark of small molecule modeling, evaluating the efficacy of both traditional machine learning and deep learning on predicting molecular properties. A recent work [70] proposes a benchmark for evaluating the performance of Large Language Models (LLMs) on molecular property prediction. In the field of protein modeling, the renowned CASP [32] competition is dedicated to establishing standards for protein structure prediction. Also, benchmark datasets are constructed for machine learning guided protein engineering [46, 19], protein design [20] and protein function annotation [63, 69]. Benchmark datasets are also established for other biomolecules like DNAs [28, 41] and RNAs [60]. In this work, we take the initiative of building a glycan machine learning benchmark for comprehensive glycan understanding.
@@ 3.1 Glycan Taxonomy Prediction
Scientific significance. The taxonomy of glycans lays the foundation for glycomics research [4]. Biologists commonly classify glycans based on their origin under the hierarchical system of domain, kingdom, phylum, class, order, family, genus and species. Such a systematic classification helps us compare the similarities and differences between glycans, which further facilitates the study of glycan structures and functions. Also, the glycan taxonomy helps us understand the process of biological evolution. By comparing the structures of glycans in different organisms, we can infer their phylogenetic relationships and possible changes that may occur during evolution. Therefore, it could be very helpful to have an accurate glycan taxonomy predictor based on machine learning.
Task definition. In the GlycanML benchmark, we study glycan taxonomy prediction on domain, kingdom, phylum, class, order, family, genus and species levels, leading to eight individual tasks. These tasks are formulated as classification problems with 4, 11, 39, 101, 210, 415, 922 and 1,737 biological categories, respectively. Taking the class imbalance into consideration, we report Macro-F1 score for each task.
@@ 3.2 Glycan Immunogenicity Prediction
Scientific significance. Predicting the immunogenicity of glycans is of great significance for vaccine design and disease treatment. (1) Glycans are key components in many vaccines, especially in bacterial vaccines. By predicting the immunogenicity of glycans, researchers can design more effective vaccine formulations, thereby improving the protective effect of vaccines [29]. (2) In addition, certain glycans can inhibit tumor growth by activating the immune system [3], and therefore accurately predicting glycan immunogenicity can help optimize tumor treatment strategies.
@@ 4.1 Representations
Model Input Layer Hidden Layers Output Layer #Params. Sequence encoders Shallow CNN [49] 128-dim. token embedding 2 $\times$ 1D conv. layers (hidden dim.: 128; max pooling over all tokens 191.7K kernel size: 5; stride: 1; padding: 2) ResNet [24] 512-dim. token embedding 8 $\times$ residual blocks (hidden dim.: 512; attentive weighted sum over all tokens 11.4M + 512-dim. positional embedding kernel size: 3; stride: 1; padding: 1) LSTM [26] 640-dim. token embedding 3 $\times$ bidirectional LSTM layers weighted sum over all tokens 26.7M (hidden dim.: 640) + linear (output dim.: 640) + Tanh Transformer [56] 512-dim. token embedding 4 $\times$ Transformer blocks (hidden dim.: 512; linear (output dim.: 512) + Tanh 21.4M + 512-dim. positional embedding #attn. heads: 8; activation: GELU) upon [CLS] token Homogeneous GNNs GCN [31] 128-dim. node embedding 3 $\times$ GCN layers concat. mean & max pooling 67.8K GAT [57] 128-dim. node embedding 3 $\times$ GAT layers (#attn. heads: 2) concat. mean & max pooling 69.4K GIN [62] 128-dim. node embedding 3 $\times$ GIN layers concat. mean & max pooling 117.4K Heterogeneous GNNs MPNN [22] 128-dim. node & edge embedding 3 $\times$ MPNN layers Set2Set pooling [59] (#steps: 3) 4.0M RGCN [48] 128-dim. node embedding 3 $\times$ RGCN layers concat. mean & max pooling 4.2M CompGCN [55] 128-dim. node embedding 3 $\times$ CompGCN layers concat. mean & max pooling 150.4K All-Atom Molecular Encoders Graphormer [66] 768-dim. node embedding 8 $\times$ Graphormer layers concat. mean & max pooling 28.5M GraphGPS [45] 512-dim. node & edge embedding 6 $\times$ GraphGPS layers concat. mean & max pooling 17.4M
@@ 4.2 Baselines
We include four types of models in our benchmark, i.e., sequence encoders for modeling glycan tokenized sequences, homogeneous and heterogeneous GNNs for modeling glycan planar graphs, and all-atom molecular encoders for modeling the all-atom molecular graphs of glycans, with 12 baseline models in total. Their detailed architectures are provided in Table 2.
Sequence encoders. We study the performance of four typical sequence encoders. Inspired by the success of shallow CNNs in modeling biological sequences like protein sequences [49, 63], we investigate (1) a 2-layer shallow CNN model along with (2) a deep residual network (ResNet) [24] with 8 hidden layers. These two CNN models mainly focus on capturing local information in glycan sequences. To investigate the importance of long context modeling for glycan understanding, we also include (3) a 3-layer bidirectional LSTM [26] and (4) a 4-layer Transformer encoder [56].
@@
All-atom molecular encoders. Essentially, glycans are a kind of macromolecules with hundreds of atoms, and therefore we can directly employ all-atom molecular encoders designed for small molecules to model glycans. Among existing small molecule encoders, Graphormer [66] and GraphGPS [45] achieve prominent performance, and we include them in the proposed benchmark. We follow the default architecture of these two models in their original work, as shown in Table 2.
@@ 4.3 Model Pipelines
Protein-glycan interaction prediction. Because of the additional input of protein, the protein-glycan interaction prediction task requires a different pipeline. Given a protein and a glycan, we first extract the protein representation vector with a protein encoder (e.g., the ESM-1b pre-trained protein language model [47] in this work) and extract the glycan representation vector with a glycan encoder, and these two vectors are then concatenated and sent to an MLP head for interaction prediction.
@@ 4.4 Multi-Task Learning
On this benchmark, we study 6 representative MTL methods that focus on loss design and model optimization under the MTL setting. All these methods use the network architecture with hard parameter sharing [68], where all tasks share a common backbone encoder, and each task owns its individual prediction head. We introduce these methods below, with an abbreviation after each one.
@@
  • Gradient Normalization (GN) [15]: However, regarding all tasks equally is suboptimal, considering the varying difficulties of different tasks. Therefore, this method employs a weighted loss summation $\mathcal{L}_{\mathrm{GN}}=\sum_{i=1}^{8}w_{i}\mathcal{L}_{i}$, where the weights satisfy: $\sum_{i=1}^{8}w_{i}=8$. The main idea of gradient normalization is that different tasks should be trained at similar rates (i.e., similar speed of convergence). To achieve this goal, authors first deem the L2 norm of per-task gradient as the training rate of the task: $r_{i}=||\nabla_{\theta}w_{i}\mathcal{L}_{i}||_{2}$ ($\theta$ denotes model parameters), and all tasks are then pushed to have similar training rates by optimizing the loss $\mathcal{L}(w_{1},\cdots,w_{8})=\sum_{i=1}^{8}||r_{i}-\bar{r}||_{1}$ ($\bar{r}=(\sum_{i=1}^{8}r_{i})/8$). For each training step, this loss is first optimized w.r.t. loss weights $\{w_{i}\}_{i=1}^{8}$, and, using the updated loss weights, the MTL loss $\mathcal{L}_{\mathrm{GN}}$ optimizes whole model parameters.
  • Temperature Scaling (TS) [30]: For classification tasks, the sharpness of categorical distribution represents prediction uncertainty, further implying task difficulty. Inspired by this fact, the TS method seeks to weigh different tasks by scaling their classification logits. In this way, each task loss is defined as $\mathcal{L}^{\mathrm{TS}}_{i}=-\log(\mathrm{Softmax}(f_{\theta}(y|x)/\sigma_{i}^{2}))$ ($i=1,\cdots,8$), where $f_{\theta}(y|x)$ represents the classification logit of sample $x$ at class $y$, and $\sigma_{i}$ denotes the task-specific temperature parameter for scaling. The temperature parameters are learned along with the whole model.
@@ 5.1 Experimental Setups
Model setups. For glycan taxonomy, immunogenicity and glycosylation type prediction tasks, upon the glycan embedding extracted by the glycan encoder, we use an MLP with 2 hidden layers and a ReLU nonlinearity in between to perform prediction. For protein-glycan interaction prediction, we use the ESM-1b pre-trained protein language model [47] to extract protein embedding, and, upon the concatenation of protein and glycan embeddings, the binding affinity is predicted by a 2-layer MLP with ReLU activation.
@@ 5.2 Benchmark Results on Single-Task Learning
In Table 3, we report the single-task performance of 12 representative glycan encoders on benchmark tasks. We measure the comprehensive performance of a model with its weighted mean rank over all tasks, where each taxonomy prediction task owns the weight of $1/8$ and each of the other three tasks owns the weight of 1, so as to balance between different types of tasks. Based on these results, we highlight the following findings:
@@
  • A simple shallow CNN is surprisingly effective. It is observed that the 2-layer shallow CNN ranks among the top three on 5 out of 11 tasks, and it ranks fifth place in terms of weighted mean rank. We thereby demonstrate that such a shallow CNN model is sufficient to produce informative glycan representations and lead to competitive performance, which aligns with previous findings that shallow CNNs can well model biological sequences like protein sequences [49, 63].
@@
Method Domain Kingdom Phylum Class Order Family Genus Species Mean Macro-F1 Backbone encoder: Shallow CNN Single-Task 0.629(0.005) 0.559(0.024) 0.388(0.024) 0.342(0.020) 0.238(0.016) 0.200(0.014) 0.149(0.009) 0.115(0.008) 0.327(0.003) N-MTL 0.619(0.009) 0.549(0.040) 0.333(0.029) 0.344(0.015) 0.213(0.015) 0.167(0.014) 0.150(0.008) 0.121(0.010) 0.312(0.010) GN 0.646(0.043) 0.497(0.045) 0.344(0.064) 0.312(0.002) 0.233(0.042) 0.206(0.027) 0.175(0.038) 0.137(0.038) 0.319(0.012) TS 0.642(0.031) 0.609(0.059) 0.393(0.061) 0.355(0.013) 0.249(0.019) 0.187(0.009) 0.141(0.010) 0.113(0.009) 0.336(0.014) UW 0.632(0.005) 0.620(0.046) 0.409(0.025) 0.320(0.011) 0.230(0.012) 0.179(0.012) 0.152(0.001) 0.118(0.008) 0.332(0.005) DWA 0.619(0.012) 0.536(0.018) 0.354(0.013) 0.330(0.041) 0.218(0.011) 0.152(0.018) 0.138(0.015) 0.099(0.012) 0.305(0.013) DTP 0.633(0.004) 0.556(0.023) 0.341(0.041) 0.303(0.028) 0.213(0.034) 0.171(0.010) 0.145(0.011) 0.107(0.006) 0.309(0.017) Backbone encoder: RGCN Single-Task 0.633(0.001) 0.647(0.054) 0.462(0.033) 0.373(0.036) 0.251(0.012) 0.203(0.008) 0.164(0.003) 0.146(0.004) 0.360(0.009) N-MTL 0.641(0.001) 0.646(0.066) 0.402(0.042) 0.376(0.054) 0.231(0.025) 0.190(0.007) 0.154(0.018) 0.119(0.007) 0.345(0.020) GN 0.626(0.002) 0.612(0.050) 0.379(0.011) 0.364(0.017) 0.233(0.010) 0.169(0.006) 0.148(0.009) 0.119(0.006) 0.331(0.005) TS 0.640(0.003) 0.652(0.050) 0.461(0.013) 0.388(0.031) 0.267(0.024) 0.209(0.010) 0.174(0.002) 0.151(0.002) 0.368(0.016) UW 0.362(0.014) 0.568(0.045) 0.395(0.042) 0.386(0.045) 0.245(0.015) 0.199(0.009) 0.165(0.005) 0.133(0.012) 0.341(0.019) DWA 0.587(0.095) 0.616(0.112) 0.400(0.035) 0.379(0.024) 0.230(0.017) 0.177(0.023) 0.151(0.015) 0.110(0.005) 0.331(0.032) DTP 0.647(0.019) 0.637(0.083) 0.399(0.069) 0.370(0.023) 0.266(0.031) 0.228(0.039) 0.190(0.043) 0.158(0.035) 0.362(0.007)

### v3

@@ Abstract
Glycans are basic biomolecules and perform essential functions within living organisms. The rapid increase of functional glycan data provides a good opportunity for machine learning solutions to glycan understanding. However, there still lacks a standard machine learning benchmark for glycan property and function prediction. In this work, we fill this blank by building a comprehensive benchmark for Glycan Machine Learning (GlycanML). The GlycanML benchmark consists of diverse types of tasks including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. Glycans can be represented by both sequences and graphs in GlycanML, which enables us to extensively evaluate sequence-based models and graph neural networks (GNNs) on benchmark tasks. Furthermore, by concurrently performing eight glycan taxonomy prediction tasks, we introduce the GlycanML-MTL testbed for multi-task learning (MTL) algorithms. Also, we evaluate how taxonomy prediction can boost other three function prediction tasks by MTL. Experimental results show the superiority of modeling glycans with multi-relational GNNs, and suitable MTL methods can further boost model performance. We provide all datasets and source codes at https://github.com/GlycanML/GlycanML and maintain a leaderboard at https://GlycanML.github.io/project.
@@ 1 Introduction
There are some existing works that employ machine learning models to predict the species origins of glycans [11, 12], glycosylation phenomenon [40, 34] and the ability of glycans to induce immune response [58, 38]. These works mainly aim to solve one or several related glycan understanding problems. However, there still lacks a comprehensive benchmark studying the general effectiveness of various machine learning models on predicting diverse glycan properties and functions, which hinders the progress of machine learning for glycan understanding. As a matter of fact, comprehensive benchmark studies greatly facilitate the machine learning research of other biomolecules like small molecules [59, 51], proteins [44, 61] and nucleic acids [57, 39].
@@
Therefore, in this work, we take the initiative of building a Glycan Machine Learning (GlycanML) benchmark featured with diverse types of tasks and multiple glycan representation structures. The GlycanML benchmark consists of 11 benchmark tasks for understanding important glycan properties and functions, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. For each task, we carefully split the benchmark dataset to evaluate the generalization ability of machine learning models in real-world scenarios. For example, in glycan taxonomy prediction, we leave out the glycans with unseen structural motifs during training for validation and test, which simulates the classification of newly discovered glycans in nature with novel molecular structures.
@@
The GlycanML benchmark accommodates two glycan representation structures, i.e., glycan tokenized sequences and glycan planar graphs. For each structure, we adopt suitable machine learning models for representation learning, where sequence encoders such as CNN [23], LSTM [25] and Transformer [53] are employed for glycan sequence encoding, and both homogeneous GNNs [29, 54, 60] and heterogeneous GNNs [21, 46, 52] are used to encode glycan graphs. In addition, two state-of-the-art small molecule encoders [64, 43] are also studied for all-atom glycan modeling. We evaluate each model on all benchmark tasks to study its general effectiveness.
The GlycanML benchmark also provides a testbed, namely GlycanML-MTL, for multi-task learning (MTL) algorithms, where an MTL method is asked to simultaneously solve eight glycan taxonomy prediction problems which are highly correlated. The performance on this testbed measures how well an MTL method can transfer the knowledge learned from different glycan taxonomies, e.g., transferring between species-level classification and genus-level classification. Taking a step further, we study the effect of MTL on boosting three function prediction tasks (i.e., immunogenicity, glycosylation and interaction prediction) by concurrently performing taxonomy prediction.
@@
Benchmark results show that the RGCN model [46], a typical heterogeneous GNN, performs best on most benchmark tasks, and a simple two-layer CNN can surprisingly achieve competitive performance by using only condensed sequential information of glycan structures. The MTL methods with elaborate task-reweighting strategies can further enhance the performance of taxonomy prediction, and learning the immunogenicity and glycosylation type of glycans jointly with their taxonomies also achieves benefits, showing the potential of MTL on boosting glycan understanding. We hope the GlycanML benchmark will spark the interest of studying glycoscience with machine learning.
@@ 2 Related Work
Glycan machine learning. With the expanding size of experimental glycomics datasets, the integration of machine learning techniques into glycoinformatics shows considerable promise [8, 34]. Early approaches use traditional machine learning algorithms (e.g., SVMs) to learn patterns from mass spectrometry data [31, 35], predict glycosylation sites [13, 33, 42], and classify glycans [62]. Recently, thanks to the advancements in deep learning and new glycomics datasets, there has been a rise in studies applying deep learning to glycan and glycosylation modeling. These works seek to identify N-glycosylated sequon [40], model glycan 3D structures [5, 16] and predict various glycan properties and functions [9, 12, 18, 38, 14, 1].
@@
Biological machine learning benchmarks. To evaluate the performance of different machine learning methods in modeling biomolecules, it is necessary to establish large-scale standardized benchmarks. MoleculeNet [59] is a widely-used benchmark of small molecule modeling, evaluating the efficacy of both traditional machine learning and deep learning on predicting molecular properties. In the field of protein modeling, the renowned CASP [30] competition is dedicated to establishing standards for protein structure prediction. Also, benchmark datasets are constructed for machine learning guided protein engineering [44, 19], protein design [20] and protein function annotation [61, 67]. Benchmark datasets are also established for other biomolecules like DNAs [26, 39] and RNAs [57]. In this work, we take the initiative of building a glycan machine learning benchmark for comprehensive glycan understanding.
@@ 3.1 Glycan Taxonomy Prediction
Scientific significance. The taxonomy of glycans lays the foundation for glycomics research [4]. Biologists commonly classify glycans based on their origin under the hierarchical system of domain, kingdom, phylum, class, order, family, genus and species. Such a systematic classification helps us compare the similarities and differences between glycans, which further facilitates the study of glycan structures and functions. Also, the glycan taxonomy helps us understand the process of biological evolution. By comparing the structures of glycans in different organisms, we can infer their phylogenetic relationships and possible changes that may occur during evolution. Therefore, it is very helpful to have an accurate glycan taxonomy predictor based on machine learning.
Task definition. We study glycan taxonomy prediction on domain, kingdom, phylum, class, order, family, genus and species levels, leading to eight individual tasks. These tasks are formulated as classification problems with 4, 11, 39, 101, 210, 415, 922 and 1,737 biological categories, respectively. Taking the class imbalance into consideration, we report Macro-F1 score for each task.
@@ 3.2 Glycan Immunogenicity Prediction
Scientific significance. Predicting the immunogenicity of glycans is of great significance for vaccine design and disease treatment. (1) Glycans are key components in many vaccines, especially in bacterial vaccines [27]. By predicting the immunogenicity of glycans, researchers can design more effective vaccine formulations. (2) In addition, certain glycans can inhibit tumor growth by activating the immune system [3], and therefore accurately predicting glycan immunogenicity can help optimize tumor treatment strategies.
@@ 4.2 Baselines
We include four types of models in our benchmark, i.e., sequence encoders for modeling glycan tokenized sequences, homogeneous and heterogeneous GNNs for modeling glycan planar graphs, and all-atom molecular encoders for modeling the all-atom molecular graphs of glycans, with 12 baseline models in total. Their detailed architectures are provided in Appendix A.
Sequence encoders. We study the performance of four typical sequence encoders. Inspired by the success of shallow CNNs in modeling biological sequences like protein sequences [47, 61], we investigate (1) a 2-layer shallow CNN along with (2) a deep residual network (ResNet) [23] with 8 hidden layers. These two CNN models mainly focus on capturing local information in glycan sequences. To investigate the importance of long context modeling for glycan understanding, we also include (3) a 3-layer bidirectional LSTM [25] and (4) a 4-layer Transformer encoder [53].
@@
All-atom molecular encoders. Essentially, glycans are a kind of macromolecules with hundreds of atoms, and thus we can directly use all-atom molecular encoders for small molecules to model glycans. Among existing small molecule encoders, Graphormer [64] and GraphGPS [43] achieve prominent performance, and we include them in this benchmark.
@@ 4.3 Model Pipelines
Protein-glycan interaction prediction. Because of the additional input of protein, protein-glycan interaction prediction requires a different pipeline. Given a protein and a glycan, we first extract the protein representation with a protein encoder (e.g., the ESM-1b protein language model [45] used in this work) and extract the glycan representation with a glycan encoder, and these two representations are then concatenated and sent to an MLP head for interaction prediction.
@@ 4.4 Multi-Task Learning
On this benchmark, we study 6 representative MTL methods for MTL loss design and MTL model optimization. All these methods use the network architecture with hard parameter sharing [66], where all tasks share a common backbone encoder, and each task owns its individual prediction head. We introduce these methods below, with an abbreviation after each one.
@@
  • Gradient Normalization (GN) [15]: However, regarding all tasks equally is suboptimal, considering the varying difficulties of different tasks. Therefore, this method employs a weighted loss summation $\mathcal{L}_{\mathrm{GN}}=\sum_{i=1}^{8}w_{i}\mathcal{L}_{i}$, where the weights satisfy: $\sum_{i=1}^{8}w_{i}=8$. The main idea of gradient normalization is that different tasks should be trained at similar rates (i.e., similar speed of convergence). To achieve this goal, authors first deem the L2 norm of per-task gradient as the training rate of the task: $r_{i}=||\nabla_{\theta}w_{i}\mathcal{L}_{i}||_{2}$ ($\theta$ denotes model parameters), and all tasks are then pushed to have similar training rates by optimizing the loss $\mathcal{L}(w_{1},\cdots,w_{8})=\sum_{i=1}^{8}||r_{i}-\bar{r}||_{1}$ ($\bar{r}=(\sum_{i=1}^{8}r_{i})/8$). For each training step, this loss is first optimized w.r.t. loss weights $\{w_{i}\}_{i=1}^{8}$, and, using the updated loss weights, the whole model is optimized by $\mathcal{L}_{\mathrm{GN}}$.
  • Temperature Scaling (TS) [28]: For classification tasks, the sharpness of categorical distribution represents prediction uncertainty, further implying task difficulty. Inspired by this fact, the TS method weighs different tasks by scaling their classification logits. In this way, each task loss is defined as $\mathcal{L}^{\mathrm{TS}}_{i}=-\log(\mathrm{Softmax}(f_{\theta}(y|x)/\sigma_{i}^{2}))$ ($i=1,\cdots,8$), where $f_{\theta}(y|x)$ is the classification logit of sample $x$ at class $y$, and $\sigma_{i}$ denotes the task-specific temperature parameter for scaling. The temperature parameters are learned along with the whole model.
@@ 5.1 Experimental Setups
Model setups. For glycan taxonomy, immunogenicity and glycosylation type prediction, upon the glycan representation extracted by the glycan encoder, we use a 2-layer MLP with ReLU activation to perform prediction. For protein-glycan interaction prediction, we use the ESM-1b protein language model [45] to extract protein representation, and, upon the concatenation of protein and glycan representations, the binding affinity is predicted by a 2-layer MLP with ReLU activation.
@@ 5.2 Benchmark Results on Single-Task Learning
In Table 2, we report the single-task performance of 12 representative glycan encoders. We measure the comprehensive performance of a model with its weighted mean rank over all tasks, where each taxonomy prediction task weighs $1/8$ and each of the other three tasks weighs 1, so as to balance between different types of tasks. Based on these results, we highlight the following findings:
@@
  • A simple shallow CNN is surprisingly effective. It is observed that the 2-layer shallow CNN ranks among the top three on 5 out of 11 tasks, and it ranks fifth place in terms of weighted mean rank. Therefore, such a shallow CNN model is sufficient to produce informative glycan representations and achieve competitive performance, aligning with previous findings that shallow CNNs can well model biological sequences [47, 61].